OpenStudy (anonymous):

The ratio 25:45 in simplest form is 5:9 5:8 9:5 Can anyone help???

OpenStudy (uri):

5:9

OpenStudy (anonymous):

\[\frac{ 25 }{ 45 }=\frac{ 5*5 }{ 5*9 }=\frac{ 5 }{ 9 }\]

OpenStudy (anonymous):

Thank You :D

OpenStudy (anonymous):

yw